OpenStudy (anonymous):

5a^(2)+16a+3 step by step work shown

OpenStudy (anonymous):

(5a + 1)*(a +3) = 5a^2 + 15a + a + 3 = original equation

OpenStudy (anonymous):

to understand this problem, you need to practice more and recognize the problem at the first look you have 5 (from 5a^2) and 16 (from 16a) and 3 to have 3 we need 1*3 = 3 that's why we have (x +1)*(y+3) now to have 5a^2 we need 5a*a = 5a^2 that's why we have (5a+1)*(a+3) the 16a will be automatically added up